# Benchmarks and Challenges in Pose Estimation for Egocentric Hand Interactions with Objects

### v1

@@ 5.1 Results
Here we benchmark results of valid submissions for state-of-the-art comparison in AssemblyHands and ARCTIC as well as other more recent baselines. In particular, for AssemblyHands, we report egocentric hand pose estimation results. For ARCTIC, we report results for the allocentric and egocentric test sets.
AssemblyHands benchmark: Table 3 shows the final test scores on the AssemblyHands dataset. The methods listed in the table surpass the baseline (Base) with a test score of 20.69 MPJPE. Notably, the methods JHands and PICO-AI achieve a nearly 40 % reduction over the baseline. The methods FRDC and Phi-AI improve the test score by 20.3 % and 16.5 % against the baseline, respectively.
ARCTIC benchmark: Table 4 presents the comparative performance of methods in the ARCTIC dataset, where ArcticNet-SF serves as the initial benchmark. The majority surpass ArcticNet-SF in both allocentric and egocentric views, except for UVHand, which underperforms due to incomplete training. In the egocentric perspective, AmbiguousHands excels in creating smooth, consistent mesh motions (refer to MDev and ACCh/o metrics). Notably, JointTransformer stands out by significantly lowering CDev errors by $32.7\%$ in allocentric and $27.2\%$ in egocentric settings compared to the baseline.
@@ 5.2 AssemblyHandsanalysis
Figure 3: Effect of distortion in AssemblyHands. The images officially released in the dataset (left) include highly stretched areas near the image edge (original crop). The method JHands proposes a perspective crop based on a virtual camera set on the cropping area, allowing to correct the distortion.
Table 5: Results of multi-view fusion in AssemblyHands. We use the Base result to show performance before and after fusion. Missing instances per view are denoted as “Miss(%)”.
@@ 5.3 ARCTICanalysis
Figure 4: Performance comparison between egocentric and allocentric. (a) The number of times harder the egocentric view is to allocentric (see text for details). (b) Egocentric view performance for each method across objects. (c) Allocentric view performance for each method across objects. Best viewed in color.
Egocentric-allocentric comparison: Figure 4a compares the performance between egocentric and allocentric views. In particular, we compute a ratio between the metric values of the egocentric and allocentric view to indicates how many times the egocentric view is more difficult than the allocentric. Since success rate is a metric whose value is positively correlated to performance, we take its reciprocal ratio. We average the ratios across methods and actions.
We observe that hand-pose related metrics such as MPJPE and ACCh are less than 1.0 on average (see blue color cells), meaning the the egocentric view is easier than allocentric view. This is because most allocentric cameras in ARCTIC are meters away from the subject while the egocentric camera is often close-up, offering higher hand visibility. Relative translation metrics between hand and object such as MRRPErl are also easier in the egocentric view because estimating translation is more difficult from further cameras.
Object reconstruction performance faces unique challenges, as indicated by the red color cells. In the egocentric view, reconstructing accurate object surfaces, articulation (AAE), and hand-object contact (CDev) is notably more difficult. This increased challenge stems from objects being frequently positioned at the image edges and obscured by human arms in the egocentric perspective. Additionally, object poses exhibit greater diversity in the egocentric view compared to the allocentric view due to varying camera angles and occlusions. In detail, while a static camera uses consistent camera extrinsics across a sequence, an egocentric camera’s extrinsics change with each frame, leading to a higher diversity in the camera-view object 6D poses. This diversity makes object pose estimation more challenging in egocentric views. Figure 5 illustrates these challenges using the best performing method JointTransformer. Despite reasonable hand poses, object poses are significantly affected by occlusions from hands and arms, and the egocentric view undergoes considerable changes in a sequence.
@@
Object-wise evaluation: Figure 4b and Figure 4c shows a breakdown of method performance on different objects in the egocentric and allocentric view test sets. In both cases, the best-performing method is JointTransformer. The hardest object to reconstruct hand and object meshes with good contact consistency (measured by CDev) is the microwave in both settings. In particular, objects are more difficult to estimate in the egocentric view than in allocentric views, indicated by the higher errors for all methods.
Action-wise evaluation: Figure 7 compares performance of different methods in “grab” and “use” actions. In ARCTIC, there are sequences to interact with the object with two types of actions by either not articulating the object, or allowing object articulation. Interestingly, the “grab” motion is more challenging in both egocentric and allocentric views. We hypothesize that this is because there are more diverse object poses for “grab” motions because during object articulation the participants often focus on articulation instead of object manipulation.
Effect of model size: Figure 7 shows the effect of model size on hand-object contact performance (measured by CDev) for reconstruction on the allocentric validation set. We see that most methods use the ResNet50 , among which the best performing method is JointTransformer. We also show that as we scale up the backbone trainable parameters, JointTransformer consistently decreases the CDev error. Note that the x-axis is in log-scale. The model performance of JointTransformer was 30.530.5mm for ViT-L and 29.029.0mm for ViT-G (almost 10 times more parameters than ViT-L). Interestingly, while JointTransformer leverages large-scale backbone ViT-L and ViT-G, the backbone weights are frozen, yet yielding the best results. This leads to a direction to leverage these large-scale foundational backbones for hand-object reconstruction.
@@ 6 Conclusion
In this paper, we introduce the HANDS23 challenge and provide analysis based on the results of the top submitted methods and more recent baselines on the leaderboards. We organize and compare submitted methods and their implementation details in terms of learning methods, architecture, pre- and post-processing techniques, and training configurations. We undertake thorough analyses on various aspects, such as with hand-object occlusions, action and object-wise evaluation, distortion correction, multi-view fusion, egocentric-allocentric comparison, and performance gain of large transformer models.
In the future, we would like to perfect the details of the benchmarks, collect more state-of-the-art methods for analysis and provide more insights for egocentric hand interactions with objects. We believe important future work to be extended from this paper includes but is not limited to efficient training using multi-view egocentric cameras, in-hand object reconstruction without prior 3D object templates, motion and temporal modeling beyond per-frame estimation, featuring more diverse egocentric interaction scenarios, recognizing actions through captured hand poses, learning robotic grasping from reconstructed hand-object pose sequences, and so forth.
Acknowledgement
This work is partially supported by the MSIT (Ministry of Science and ICT), Korea, under the ITRC support program (IITP-2024-2020-0-01789) supervised by the IITP. UTokyo is supported by JST ACT-X Grant Number JPMJAX2007, JSPS KAKENHI Grant Number JP22KJ0999, JST Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE), Grant Number JPMJAP2303. NUS is supported by the Ministry of Education, Singapore, under its MOE Academic Research Fund Tier 2 (STEM RIE2025 MOE-T2EP20220-0015).
@@ A Additional results of AssemblyHands
Figure b: Additional results of multi-view fusion in AssemblyHands. We analyze the availability of samples and performance per camera view. The bottom two cameras (cam3, cam4) out of the four cameras enable capturing hands for most of the time (>>93 % of samples). In contrast, the images from cam1 and cam2 are fewer and the error varies in different sequences.

### v2

@@ 5.1 Results
Here, we benchmark results of valid submissions for state-of-the-art comparison in AssemblyHands and ARCTIC and other more recent baselines. In particular, for AssemblyHands, we report egocentric hand pose estimation results. For ARCTIC, we report results for the allocentric and egocentric test sets.
AssemblyHands benchmark: Table 3 shows the final test scores on the AssemblyHands dataset. The methods in the table exceed the baseline (Base) with a test score of 20.69 MPJPE. Notably, the methods JHands and PICO-AI achieve a nearly 40 % reduction over the baseline. The methods FRDC and Phi-AI improve the test score by 20.3 % and 16.5 % against the baseline, respectively.
ARCTIC benchmark: Table 4 presents the comparative performance of methods in the ARCTIC dataset, where ArcticNet-SF serves as the initial benchmark. The majority surpass ArcticNet-SF in both allocentric and egocentric views, except for UVHand, which underperforms due to incomplete training. In the egocentric view, AmbiguousHands excels in creating smooth, consistent mesh motions (refer to MDev and ACCh/o metrics). Notably, JointTransformer stands out by significantly lowering CDev errors by $32.7\%$ in allocentric and $27.2\%$ in egocentric settings compared to the baseline.
@@ 5.2 AssemblyHandsanalysis
Figure 3: Effect of distortion in AssemblyHands. The officially released images in the dataset have highly stretched areas near the edges (original crop). The method JHands uses a perspective crop with a virtual camera to correct this distortion.
Table 5: Multi-view fusion in AssemblyHands. We use the Base result to show performance before and after fusion. Missing instances per view are denoted as “Miss(%)”.
@@ 5.3 ARCTICanalysis
Figure 4: Performance comparison: Egocentric vs Allocentric. (a) Comparative difficulty ratio of egocentric to allocentric views. (b) Egocentric view performance by method across objects. (c) Allocentric view performance by method across objects.
Egocentric-allocentric comparison: Figure 4a compares the performance between egocentric and allocentric views. In particular, we compute a ratio between the metric values of the egocentric and allocentric view to measure the extent of difficulty for the egocentric view compared to the allocentric view. Since success rate is a metric whose value is positively correlated to performance, we take its reciprocal ratio. We average the ratios across methods and actions.
We observe that hand pose-related metrics such as MPJPE and ACCh are less than 1.0 on average (see blue color cells), meaning the the egocentric view is easier than allocentric view. This is because most allocentric cameras in ARCTIC are meters away from the subject while the egocentric camera is often close-up, offering higher hand visibility. Relative translation metrics between hand and object such as MRRPErl are also easier in the egocentric view because estimating translation is more difficult from further cameras.
Object reconstruction performance faces unique challenges, as highlighted by the red cells. In the egocentric view, it is notably more difficult to reconstruct accurate object surfaces, articulation (AAE), and hand-object contact (CDev). This increased difficulty arises because objects are often positioned at the image edges and obscured by human arms. Additionally, object poses exhibit greater diversity in the egocentric view due to varying camera angles and occlusions. While a static camera maintains consistent camera extrinsics across a sequence, an egocentric camera’s extrinsics change with each frame, resulting in higher diversity in camera-view object 6D poses. This diversity complicates object pose estimation in egocentric views. Figure 5 illustrates these challenges using the best-performing method, JointTransformer. Despite achieving reasonable hand poses, object poses are significantly impacted by occlusions from hands and arms, and the egocentric view undergoes substantial changes throughout a sequence.
@@
Object-wise evaluation: Figure 4b and 4c break down performances on different objects in the egocentric and allocentric view test sets. The best method in both cases is JointTransformer. The hardest object to reconstruct with good contact consistency (see CDev) is the microwave in both settings due to global rotation sensitivity (see Fig. 5), though this can be mitigated by a keypoint-based approach [23]. Estimating objects in the egocentric view is more difficult than in the allocentric view, which is indicated by higher errors for all methods.
Action-wise evaluation: Figure 7 compares performance of different methods in “grab” and “use” actions. In ARCTIC, there are sequences to interact with the object with two types of actions by either not articulating the object, or allowing object articulation. Interestingly, the “grab” motion is more challenging in egocentric and allocentric views. We hypothesize that this is because there are more diverse object poses for “grab” motions since during object articulation, the participants often focus on articulation instead of object manipulation.
Effect of model size: Figure 7 illustrates the impact of model size on hand-object contact performance, measured by CDev, for reconstruction on the allocentric validation set. Most methods utilize ResNet50, with JointTransformer being the top performer. As the trainable parameters in the backbone increase, JointTransformer consistently reduces the CDev error. Note that the x-axis is in log-scale. JointTransformer achieved a CDev error of 30.530.5mm with ViT-L and 29.029.0mm with ViT-G, which has ten times more parameters than ViT-L. Interestingly, JointTransformer uses frozen weights in the large-scale ViT-L and ViT-G backbones, yet achieves the best results. This suggests a potential direction for leveraging large-scale foundational backbones for hand-object reconstruction.
@@ 6 Conclusion
In this paper, we introduce the HANDS23 challenge and provide analysis based on the results of the top submitted methods and more recent baselines on the leaderboards. We organize and compare the submissions and their implementation details based on the learning methods, architecture, pre- and post-processing techniques, and training configurations. We thoroughly analyze various aspects, such as hand-object occlusions, action and object-wise evaluation, distortion correction, multi-view fusion, egocentric-allocentric comparison, and performance gain of large transformer models.
Future directions: There are several future directions that the community can take. For example, one can explore more efficient training using multi-view egocentric cameras, leveraging 3D foundation priors [54, 38] to regularize template-free hand-object reconstruction [14], estimating hand-object poses with more expressive representations (e.g., heatmap-based approaches [23]), incorporating motion and temporal modeling [17], featuring more diverse egocentric interaction scenarios, recognizing actions through captured hand poses [62], learning robotic grasping from reconstructed hand-object pose sequences, and so forth.
Acknowledgements
This work is supported by the Ministry of Science and ICT, Korea, under the ITRC program (IITP-2024-2020-0-01789). supervised by the IITP. UTokyo is supported by JST ACT-X Grant Number JPMJAX2007, JSPS KAKENHI Grant Number JP22KJ0999, JST Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE) Grant Number JPMJAP2303. NUS is supported by the Ministry of Education, Singapore, under its MOE Academic Research Fund Tier 2 (STEM RIE2025 MOE-T2EP20220-0015).
@@ A Additional results of AssemblyHands
Figure b: Additional results of multi-view fusion in AssemblyHands. We analyze the availability of samples and performance per camera view. The two lowest cameras (cam3, cam4) out of the four cameras allow us to capture hands most of the time (>>93 % of samples). In contrast, the images from cam1 and cam2 are fewer and the error varies in different sequences.